### v1

# A linear independence criterion for certain infinite series with polynomial orders

Abstract
Let $q$ be a Pisot or Salem number. Let $f_{j}(x)$ $(j=1,2,\dots)$ be integer-valued polynomials of degree $\geq 2$ with positive leading coefficients, and let $\{a_{j}(n)\}_{n\geq 1}$ $(j=1,2,\dots)$ be sequences of algebraic integers in the field $\mathbb{Q}(q)$ with suitable growth conditions. In this paper, we investigate linear independence over $\mathbb{Q}(q)$ of the numbers
$1,\qquad\sum_{n=1}^{\infty}\frac{a_{j}(n)}{q^{f_{j}(n)}}\quad(j=1,2,\dots).$
In particular, when $a_{j}(n)$ $(j=1,2,\dots)$ are polynomials of $n$, we give a linear independence criterion for the above numbers.
Keywords: Linear independence, Jacobi theta function, Pisot number, Salem number
Mathematics Subject Classification (2020). Primary 11J72.
1 Introduction and main results
In 1957, P. Erdős [7] proved that if $\ell\geq 1$ is an integer and $1<n_{1}<n_{2}<\cdots$ are integers with $\limsup_{k\to\infty}n_{k}/k^{\ell}=\infty$, then for any integer $b\geq 2$, the number $\sum_{k=1}^{\infty}1/b^{n_{k}}$ is transcendental or an algebraic number with degree at least $\ell+1$. For example, we see that if $f(x)$ is an integer-valued polynomial of degree $d\geq 2$ with $1<f(1)<f(2)<\cdots$, then for any integer $b\geq 2$, the number $\sum_{n=1}^{\infty}1/b^{f(n)}$ is transcendental or an algebraic number with degree at least $d$.
On the other hand, in 2019, V. Kumar [9] obtained linear independence result for certain infinite series with monomial orders. More precisely, he proved that if $k\geq 2$ and $1\leq a_{1}<a_{2}<\dots<a_{m}$ are integers with $k]{a_{i}/a_{j}}\notin\mathbb{Q}$ $(i\neq j)$, then for any integer $b\geq 2$, the numbers
$1,\quad\sum_{n=1}^{\infty}\frac{1}{b^{a_{1}n^{k}}},\quad\sum_{n=1}^{\infty}\frac{1}{b^{a_{2}n^{k}}},\dots,\sum_{n=1}^{\infty}\frac{1}{b^{a_{m}n^{k}}}$
are linearly independent over $\mathbb{Q}$. Recently, S. Murakami and Y. Tachiya [11] gave a generalization of Kumar’s result by showing that, for any integer $b\geq 2$, the set of the numbers
$1,\qquad\sum_{n=1}^{\infty}\frac{1}{b^{in^{j}}}\quad(i=1,2,\dots,\ j=2,3,\dots)$
is linearly independent over $\mathbb{Q}$.
In this paper, we study linear independence of the values of certain infinite series with polynomial orders, and as a corollary, get a generalization of linear independence result of Murkami and Tachiya [11]. Moreover, we give a linear independence criterion under specific conditions.
Before stating our results, we need some preparation. We write $f(x)\equiv g(x)$ if the polynomials $f(x)$ and $g(x)$ are identically equal, and $f(x)\not\equiv g(x)$ otherwise. Recall that the polynomial $f(x)$ is called an integer-valued polynomial if the value $f(n)$ is an integer for every positive integer $n$. Note that all coefficients of $f(x)$ are rational numbers if $f(x)$ is an integer-valued polynomial. In addition, a Pisot (resp. Salem) number are defined as a real algebraic integer greater than one, all of whose Galois conjugates other than itself have absolute value less than one (resp. less than or equal to one, and at least one conjugate has absolute value exactly one). Clearly, all of integers greater than one are Pisot numbers. Our results are the following.
Theorem 1.1.
Let $q$ be a Pisot or Salem number. Let $\ell\geq 1$ be an integer and $f_{j}(x)$ $(j=1,2,\dots,\ell)$ be integer-valued polynomials of degree $\geq 2$ with positive leading coefficients. Let $\{a_{j}(n)\}_{n\geq 1}$ $(j=1,2,\dots,\ell)$ be sequences of algebraic integers in the field $\mathbb{Q}(q)$ with $\log(1+|a_{j}(n)^{\sigma}|)=o(n)$ for every embedding $\sigma\colon\mathbb{Q}(q)\to\mathbb{C}$. Suppose that for any non-empty subset $\mathcal{S}$ of $\{1,2,\dots,\ell\}$, there exist integers $i\in\mathcal{S}$ and $A$ satisfying the following two conditions:
  1. (i)
    $f_{i}(x+A)\not\equiv f_{j}(Bx+C)+D$ for any integers $j\in\mathcal{S}\setminus\{i\}$, $C$, $D$ and any positive rational number $B$.
  2. (ii)
    There exists a positive integer $E$ such that $\liminf_{n\to\infty}|a_{i}(En+A)|\neq 0$.
Then, the numbers
$1,\qquad\sum_{n=1}^{\infty}\frac{a_{j}(n)}{q^{f_{j}(n)}}\quad(j=1,2,\dots,\ell)$ (1.1)
are linearly independent over $\mathbb{Q}(q)$.
We give the following corollary of Theorem 1.1.
Corollary 1.1.
Let $q$ be a Pisot or Salem number. Let $f_{j}(x)$ $(j=1,2,\dots,\ell)$ be integer-valued polynomials of degree $\geq 2$ with positive leading coefficients such that the condition (i) in Theorem 1.1 holds. Then, for any sequences $\{a_{j}(n)\}_{n\geq 1}$ $(j=1,2,\dots,\ell)$ of non-zero integers with $\log(1+|a_{j}(n)|)=o(n)$, the numbers
$1,\qquad\sum_{n=1}^{\infty}\frac{a_{j}(n)}{q^{f_{j}(n)}}\quad(j=1,2,\dots,\ell)$
are linearly independent over $\mathbb{Q}(q)$. In particular, so are the numbers
$1,\qquad\sum_{n=1}^{\infty}\frac{1}{q^{f_{j}(n)}}\quad(j=1,2,\dots,\ell).$
Now, we apply Theorem 1.1 to derive a linear independence result for the values of the Jacobi theta constants defined in $|z|<1$ by
$\theta_{2}(z)\coloneqq 2z^{\frac{1}{4}}\sum_{n=0}^{\infty}z^{n(n+1)},\quad\theta_{3}(z)\coloneqq 1+2\sum_{n=1}^{\infty}z^{n^{2}},\quad\theta_{4}(z)\coloneqq 1+2\sum_{n=1}^{\infty}(-1)^{n}z^{n^{2}}.$ (1.2)
Let $f_{1}(x)\coloneqq x(x+1)$, $f_{2}(x)\coloneqq(2x)^{2}$ and $f_{3}(x)\coloneqq(2x-1)^{2}$. Then, we immediately see that the polynomials $f_{1}(x)$, $f_{2}(x)$ and $f_{3}(x)$ satisfy the condition (i) in Theorem 1.1. Let $q$ be a Pisot or Salem number, and $\ell\geq 0$ be an integer. We define
$\alpha_{i,j}\coloneqq\sum_{n=1}^{\infty}\frac{n^{i}}{q^{f_{j}(n)}}\quad(i=0,1,\dots,\ell,\ j=1,2,3).$ (1.3)
For any algebraic integers $A_{i,j}$ $(i=0,1,\dots,\ell,\ j=1,2,3)$ in $\mathbb{Q}(q)$ not all zero, we consider the linear combination over $\mathbb{Q}(q)$ of the numbers (1.3) given by
$\sum_{j=1}^{3}\sum_{i=0}^{\ell}A_{i,j}\alpha_{i,j}=\sum_{j\in\mathcal{T}}\sum_{n=1}^{\infty}\frac{a_{j}(n)}{q^{f_{j}(n)}},$ (1.4)
where
$\mathcal{T}\coloneqq\left\{j\in\{1,2,3\}\mid\text{there exists an integer }i\ (0\leq i\leq\ell)\text{ such that }A_{i,j}\neq 0\right\}\ (\neq\emptyset)$
and $a_{j}(n)\coloneqq\sum_{i=0}^{\ell}A_{i,j}n^{i}$ $(n\geq 1)$. Clearly, the sequences $\{a_{j}(n)\}_{n\geq 1}$ $(j\in\mathcal{T})$ satisfy the conditions in Theorem 1.1. Thus, by Theorem 1.1, the numbers (1.4) does not belong to $\mathbb{Q}(q)$, namely, the numbers $1,\ \alpha_{i,j}$ $(i=0,1,\dots,\ell,\ j=1,2,3)$ are linearly independent over $\mathbb{Q}(q)$. Also, for any algebraic integers $u$ and $v$ in $\mathbb{Q}(q)$, we have
$u\sum_{n=1}^{\infty}\frac{n^{k}}{q^{n^{2}}}+v\sum_{n=1}^{\infty}\frac{(-1)^{n}n^{k}}{q^{n^{2}}}=(u+v)\sum_{n=1}^{\infty}\frac{(2n)^{k}}{q^{(2n)^{2}}}+(u-v)\sum_{n=1}^{\infty}\frac{(2n-1)^{k}}{q^{(2n-1)^{2}}}\quad(k=0,1,\dots),$
and hence we get the following example of Theorem 1.1.
Example 1.1.
Let $q$ be a Pisot or Salem number. Then, for any integer $\ell\geq 0$, the numbers
$1,\qquad\sum_{n=1}^{\infty}\frac{n^{i}}{q^{n(n+1)}},\quad\sum_{n=1}^{\infty}\frac{n^{j}}{q^{n^{2}}},\quad\sum_{n=1}^{\infty}\frac{(-1)^{n}n^{k}}{q^{n^{2}}}\quad(i,j,k=0,1,\dots,\ell)$
are linearly independent over $\mathbb{Q}(q)$. Thus, so are the values
$1,\qquad\left.\frac{d^{i}}{dz^{i}}\left(z^{-\frac{1}{4}}\theta_{2}(z)\right)\right|_{z=1/q},\quad\left.\frac{d^{j}}{dz^{j}}\theta_{3}(z)\right|_{z=1/q},\quad\left.\frac{d^{k}}{dz^{k}}\theta_{4}(z)\right|_{z=1/q}\quad(i,j,k=0,1,\dots,\ell).$
Remark 1.1.
D. Bertrand [1], and independently D. Duverney, Ke. Nishioka, Ku. Nishioka and I. Shiokawa [2] proved that for any algebraic number $\alpha$ $(0<|\alpha|<1)$ and any integer $m$ $(2\leq m\leq 4)$, the three values $\theta_{m}(\alpha)$, $\mathcal{D}\theta_{m}(\alpha)$ and $\mathcal{D}^{2}\theta_{m}(\alpha)$ are algebraically independent over $\mathbb{Q}$, where $\mathcal{D}\coloneqq z(d/dz)$ is a differential operator, by applying results of Yu. V. Nesterenko [12]. On the other hand, for any above numbers $\alpha$ and $m$, the four values $\theta_{m}(\alpha)$, $\mathcal{D}\theta_{m}(\alpha)$, $\mathcal{D}^{2}\theta_{m}(\alpha)$ and $\mathcal{D}^{3}\theta_{m}(\alpha)$ are algebraically dependent over $\mathbb{Q}$ since $\theta_{m}$ has an algebraic differential equation of the third order over $\mathbb{Q}$ (cf. [8]). In addition, Bertrand showed that for any algebraic number $\alpha$ $(0<|\alpha|<1)$, any two values among the three values $\theta_{2}(\alpha)$, $\theta_{3}(\alpha)$ and $\theta_{4}(\alpha)$ are algebraically independent over $\mathbb{Q}$. Conversely, these three values are algebraically dependent over $\mathbb{Q}$ by the Jacobi identity $\theta_{3}^{4}(z)=\theta_{2}^{4}(z)+\theta_{4}^{4}(z)$ (cf. [10]). By the way, C. Elsner and V. Kumar [6] recently derived linear independence results for functions and values related to $\theta_{3}$.
For any integer $\ell\geq 2$, we find that the polynomials
$f_{i,j}(x)\coloneqq ix^{j}\quad(i=1,2,\dots,\ell,\ j=2,3,\dots,\ell)$
satisfy the condition (i) in Theorem 1.1 (see Lemma 5.1 in Section 5). By Corollary 1.1, we get the following example.
Example 1.2.
Let $q$ be a Pisot or Salem number. Then, for any integer $\ell\geq 2$ and any sequences $\{a_{i,j}(n)\}_{n\geq 1}$ $(i=1,2,\dots,\ell,\ j=2,3,\dots,\ell)$ of non-zero integers with $\log(1+|a_{i,j}(n)|)=o(n)$, the numbers
$1,\qquad\sum_{n=1}^{\infty}\frac{a_{i,j}(n)}{q^{in^{j}}}\quad(i=1,2,\dots,\ell,\ j=2,3,\dots,\ell)$
are linearly independent over $\mathbb{Q}(q)$.
This generalizes a result of Murakami and Tachiya [11, Corollary 1.3].
Next, we give a linear independence criterion (Theorem 1.2) for the numbers (1.1) when $a_{j}(n)$ $(j=1,2,\dots,\ell)$ are polynomials of $n$. Before stating Theorem 1.2, we need some preparation.
Let $q$ be a Pisot or Salem number. The sequences $\{b_{j}(n)\}_{n\geq 0}$ $(j=1,2,\dots,\ell)$ are called linearly dependent over $\mathbb{Q}(q)$ if there exist algebraic numbers $k_{j}$ $(j=1,2,\dots,\ell)$ in $\mathbb{Q}(q)$ not all zero such that $\sum_{j=1}^{\ell}k_{j}\cdot b_{j}(n)=0$ for every integer $n\geq 0$. Otherwise, the sequences $\{b_{j}(n)\}_{n\geq 0}$ $(j=1,2,\dots,\ell)$ are called linearly independent over $\mathbb{Q}(q)$. In addition, we define an equivalence relation $\sim$ on the set of all polynomials with rational coefficients. For two polynomials $F(x)$ and $G(x)$ with rational coefficients, we write $F\sim G$ if there exist rational numbers $b>0$, $c$ and an integer $d$ such that
$F(x)\equiv G(bx+c)+d,$
and $F\not\sim G$ otherwise. Clearly, this is an equivalence relation.
Let $f_{j}(x)$ $(j=1,2,\dots,\ell)$ be polynomials with rational coefficients and $f_{j_{1}}\sim f_{j_{2}}$ for any integers $j_{1}$, $j_{2}$ $(1\leq j_{1}\leq j_{2}\leq\ell)$. Then, there exists a polynomial $g(x)$ such that there exist integers $B_{j}>0$, $C_{j}$ and $D_{j}$ $(j=1,2,\dots,\ell)$ satisfying
$g(x)\equiv f_{j}\left(\frac{x+C_{j}}{B_{j}}\right)+D_{j}\quad(j=1,2,\dots,\ell).$ (1.5)
Indeed, by $f_{1}\sim f_{j}$ for every $j=1,2,\dots,\ell$, there exist integers $r_{j}\geq 1$, $s_{j}\geq 1$, $t_{j}$ and $u_{j}$ $(j=1,2,\dots,\ell)$ such that
$f_{1}(x)\equiv f_{j}\left(\frac{s_{j}x+t_{j}}{r_{j}}\right)+u_{j}\quad(j=1,2,\dots,\ell).$
Let $K$ be the least common multiple of $s_{j}$ $(j=1,2,\dots,\ell)$ and $g(x)\coloneqq f_{1}(x/K)$. Then,
$g(x)\equiv f_{j}\left(\frac{x+t_{j}v_{j}}{r_{j}v_{j}}\right)+u_{j}\quad(j=1,2,\dots,\ell),$
where $v_{j}\coloneqq K/s_{j}$ $(j=1,2,\dots,\ell)$ are positive integers. Thus, $g(x)$ satisfies (1.5). Theorem 1.2 is the following.
Theorem 1.2.
Let $\mathcal{U}\coloneqq\bigcup_{i=1}^{m}\{(i,j)\in\mathbb{Z}_{\geq 1}^{2}\mid j=1,2,\dots,\ell_{i}\}$ and $f_{i,j}(x)$ $\bigl{(}(i,j)\in\mathcal{U}\bigr{)}$ be integer-valued polynomials of degree $\geq 2$ with positive leading coefficients such that
$\begin{cases}f_{i_{1},j_{1}}\sim f_{i_{2},j_{2}}&\text{ if }\ i_{1}=i_{2},\\ f_{i_{1},j_{1}}\not\sim f_{i_{2},j_{2}}&\text{ if }\ i_{1}\neq i_{2}.\end{cases}$
Let $g_{i}(x)$ $(i=1,2,\dots,m)$ be polynomials given by (1.5), namely, for each $i=1,2,\dots,m$, there exist integers $B_{i,j}>0$, $C_{i,j}$ and $D_{i,j}$ $(j=1,2,\dots,\ell_{i})$ satisfying
$g_{i}(x)\equiv f_{i,j}\left(\frac{x+C_{i,j}}{B_{i,j}}\right)+D_{i,j}\quad(j=1,2,\dots,\ell_{i}).$ (1.6)
Let $q$ be a Pisot or Salem number, and $P_{i,j}(x)$ $\bigl{(}(i,j)\in\mathcal{U}\bigr{)}$ be non-zero polynomials with algebraic integer coefficients in $\mathbb{Q}(q)$. Then, the numbers
$1,\qquad\sum_{n=1}^{\infty}\frac{P_{i,j}(n)}{q^{f_{i,j}(n)}}\quad\bigl{(}(i,j)\in\mathcal{U}\bigr{)}$ (1.7)
are linearly independent over $\mathbb{Q}(q)$ if and only if for each $i=1,2,\dots,m$, the sequences
[…]
are linearly independent over $\mathbb{Q}(q)$.
In Theorem 1.2, if $\deg P_{i,j}(x)$ $\bigl{(}(i,j)\in\mathcal{U}\bigr{)}$ are distinct, we immediately see that for each $i=1,2,\dots,m$, the sequences $\{\tilde{p}_{i,j}(n)\}_{n\geq 0}$ $(j=1,2,\dots,\ell_{i})$ are linearly independent over $\mathbb{Q}(q)$.
Corollary 1.2.
Let $q$ be a Pisot or Salem number. Let $f_{j}(x)$ $(j=1,2,\dots,\ell)$ be integer-valued polynomials of degree $\geq 2$ with positive leading coefficients. Let $P_{j}(x)$ $(j=1,2,\dots,\ell)$ be non-zero polynomials with algebraic integer coefficients in $\mathbb{Q}(q)$ and $\deg P_{j_{1}}(x)\neq\deg P_{j_{2}}(x)$ for every integers $j_{1},j_{2}$ $(1\leq j_{1}<j_{2}\leq\ell)$. Then, the numbers
$1,\qquad\sum_{n=1}^{\infty}\frac{P_{j}(n)}{q^{f_{j}(n)}}\quad(j=1,2,\dots,\ell)$
are linearly independent over $\mathbb{Q}(q)$.
We give another example of Theorem 1.2 in Section 5.
The present paper is organized as follows. We will prove Theorem 1.1 in Section 3, and Theorem 1.2 in Section 4. Before proving Theorem 1.1, we give Lemma 2.1 in Section 2. Lemma 2.1 is the most important part to show Theorem 1.1. Also, we will prove Theorem 1.2 by using a result of Theorem 1.1. In Section 5, we give some examples of the set of polynomials which Theorems 1.1 and 1.2 are useful.
2 Lemma
In this section, we show the following lemma which plays an important role in the proof of Theorem 1.1. Lemma 2.1 is the following.
Lemma 2.1.
Let $\ell\geq 1$ be an integer. Let $f_{j}(x)$ $(j=1,2,\dots,\ell)$ be integer-valued polynomials of degree $\geq 2$ with positive leading coefficients and $\deg f_{1}(x)\leq\deg f_{j}(x)$ $(j=1,2,\dots,\ell)$. Moreover, assume that there exists an integer $A$ such that
$f_{1}(x+A)\not\equiv f_{j}(Bx+C)+D$ (2.1)
for any integers $j\geq 2$, $C$, $D$ and any positive rational number $B$. Let $E$ be an arbitrary positive integer and let $G(x)$ be any real-valued function for $x\geq 1$ with $G(x)\to\infty$ $(x\to\infty)$ and $G(x)=o(x)$. Then, there exist infinitely many positive integers $m$ with $m\equiv A\pmod{E}$ such that
$|f_{1}(m)-f_{j}(k)|>G(m)$
for any positive integers $j$ and $k$ with $(j,k)\neq(1,m)$.
Proof.
Suppose to the contrary that there exists a positive integer $m_{0}$ such that for any positive integer $m\geq m_{0}$ with $m\equiv A\pmod{E}$, there exists a positive integer pair $(j_{m},k_{m})\neq(1,m)$ satisfying
$|f_{1}(m)-f_{j_{m}}(k_{m})|\leq G(m).$ (2.2)
In what follows, we prove that this contradicts (2.1) by following five steps below. Before Step 1, we need some preparation.
Let $d\coloneqq\deg f_{1}(x)\geq 2$ and $r_{j}/s_{j}$ $(j=1,2,\dots,\ell)$ be rational leading coefficients of $f_{j}(x)$, where $r_{j}$ and $s_{j}$ are positive coprime integers. We define
$g(x)\coloneqq\max_{\begin{subarray}{c}y\\ 1\leq y\leq x\end{subarray}}G(y)\quad(x\geq 1).$
Then, $g(x)$ $(x\geq 1)$ is an increasing function with $g(x)\to\infty$ $(x\to\infty)$ and
$g(x)=o(x).$
Also, we define
$E_{0}$ $\coloneqq Es_{1}\cdot\operatorname*{lcm}(r_{1},r_{2},\dots,r_{\ell}),$ (2.3)
$h(x)$ $\coloneqq E_{0}d\ell\bigl{(}2g(x)+1\bigr{)}.$
For every sufficiently large integer $n$, let
$\mathcal{M}_{n}\coloneqq\{m\in\mathbb{Z}_{\geq 1}\mid m\equiv A\pmod{E_{0}},\ n-h(n)<m<n\}.$ (2.4)
By the hypotheses, note that $f_{j}(x)$, $f^{\prime}_{j}(x)$ $(j=1,2,\dots,\ell)$ are increasing for large $x$. Here, for two real functions $F_{1}(x)$ and $F_{2}(x)$, we write $F_{1}(x)\asymp F_{2}(x)$ if there exist positive real numbers $c_{1}$ and $c_{2}$ independent of $x$ such that $c_{1}F_{2}(x)\leq F_{1}(x)\leq c_{2}F_{2}(x)$ for all large $x$. Also for a real function $F(x)$ and a real number $K_{n}$ dependent on an integer $n$, we write $K_{n}\asymp F(n)$ if there exist positive real numbers $c_{3}$ and $c_{4}$ independent of $n$ such that $c_{3}F(n)\leq K_{n}\leq c_{4}F(n)$ for all large integers $n$.
Step 1. For every sufficiently large integer $n$ and for each $j=2,3,\dots,\ell$, we define
$\mathcal{M}_{n}(j)\coloneqq\{m\in\mathcal{M}_{n}\mid\text{ there exists a positive integer $k_{m}$ such that }|f_{1}(m)-f_{j}(k_{m})|\leq g(n)\}.$
In this step, we show that there exists an integer $j_{0}$ $(2\leq j_{0}\leq\ell)$ such that there exist infinitely many positive integers $n$ satisfying
#$\#\mathcal{M}_{n}(j_{0})>d\bigl{(}2g(n)+1\bigr{)}.$ (2.5)
In what follows, let $n$ be every sufficiently large integer. Since $h(x)=O\bigl{(}g(x)\bigr{)}=o(x)$, we get $n-h(n)\to\infty$ $(n\to\infty)$. Then by (2.2) and (2.4), we see that for any integer $m\in\mathcal{M}_{n}$, there exists a positive integer pair $(j_{m},k_{m})\neq(1,m)$ satisfying
$|f_{1}(m)-f_{j_{m}}(k_{m})|\leq G(m)\leq\max_{\begin{subarray}{c}y\\ 1\leq y\leq n\end{subarray}}G(y)=g(n).$ (2.6)
Also, we get
$f_{1}(x_{1})\leq f_{1}(x_{2})\quad\bigl{(}1\leq x_{1}\leq n-h(n)-1\leq x_{2}\bigr{)}.$ (2.7)
In addition, since $f_{1}(x)-f_{1}(x-1)\asymp x^{d-1}$ $(d\geq 2)$, $g(x)=o(x)$ and $n-h(n)\asymp n$, we obtain
$f_{1}(x)-f_{1}(x-1)>2g(n)\quad\bigr{(}x\geq n-h(n)\bigl{)}.$ (2.8)
Then by (2.6), (2.7) and (2.8), we get $j_{m}\neq 1$ for any integer $m\in\mathcal{M}_{n}$, and hence we have $\ell\geq 2$. Thus by (2.4) and (2.6), we obtain
#$\#\bigcup_{j=2}^{\ell}\mathcal{M}_{n}(j)=\#\mathcal{M}_{n}\geq\frac{h(n)}{E_{0}}-1>d(\ell-1)\bigl{(}2g(n)+1\bigr{)}.$
Therefore, for every sufficiently large integer $n$, there exists an integer $j(n)$ $\bigl{(}2\leq j(n)\leq\ell\bigr{)}$ such that
#$\#\mathcal{M}_{n}\bigl{(}j(n)\bigr{)}>d\bigl{(}2g(n)+1\bigr{)}.$
Thus, we see that (2.5) holds.
Step 2. In what follows, let $N$ be a sufficiently large integer satisfying
#$\#\mathcal{M}_{N}(j_{0})>d\bigl{(}2g(N)+1\bigr{)}.$ (2.9)
Let $t\in\mathcal{M}_{N}(j_{0})$, $u\geq 1$, $v\geq 1$ and $w$ be integers with $t+E_{0}v\in\mathcal{M}_{N}(j_{0})$ and $u+w\geq 1$ such that
$|f_{1}(t)-f_{j_{0}}(u)|\leq g(N),$ (2.10)
$|f_{1}(t+E_{0}v)-f_{j_{0}}(u+w)|\leq g(N),$ (2.11)
and
$(t+E_{0}v)-t=\min\{|m_{1}-m_{2}|\mid m_{1},m_{2}\in\mathcal{M}_{N}(j_{0}),\ m_{1}\neq m_{2}\}.$ (2.12)
In this step, we show that $v$ and $w$ are bounded positive integers, and $\deg f_{j_{0}}(x)=\deg f_{1}(x)=d$ holds.
First if $v>2\ell$, by (2.4), (2.9) and (2.12), we obtain
$2E_{0}\ell\{d\bigl{(}2g(N)+1\bigr{)}-1\}<m_{\max}-m_{\min}<h(N)=E_{0}d\ell\bigl{(}2g(N)+1\bigr{)},$
where $m_{\max}$ and $m_{\min}$ are the maximum and minimum of integers $m\in\mathcal{M}_{N}(j_{0})$, respectively. This is a contradiction for large $N$. Thus $1\leq v\leq 2\ell$, and hence $v$ is a bounded positive integer.
Next, we show that $\deg f_{j_{0}}(x)=\deg f_{1}(x)=d$ holds and $w$ is a bounded positive integer. By (2.8), we get $f_{1}(t+E_{0}v)-f_{1}(t)>2g(N)$. Then by (2.10) and (2.11), we obtain $w\geq 1$. Also since $t,t+E_{0}v\in\mathcal{M}_{N}(j_{0})$, we get
$t,t+E_{0}v\asymp N.$ (2.13)
Let $d_{0}\coloneqq\deg f_{j_{0}}(x)\geq d$. By (2.10) and (2.11), we get
$u,u+w\asymp N^{\frac{d}{d_{0}}}.$ (2.14)
In addition, by (2.10) and (2.11), we obtain
$\left|E_{0}v\frac{f_{1}(t+E_{0}v)-f_{1}(t)}{(t+E_{0}v)-t}-w\frac{f_{j_{0}}(u+w)-f_{j_{0}}(u)}{(u+w)-u}\right|\leq 2g(N).$
By the mean value theorem, there exist positive real numbers $\alpha$ and $\beta$ with $t<\alpha<t+E_{0}v$ and $u<\beta<u+w$ such that
$|E_{0}vf^{\prime}_{1}(\alpha)-wf^{\prime}_{j_{0}}(\beta)|\leq 2g(N).$ (2.15)
By (2.13) and (2.14), we obtain
$f^{\prime}_{1}(\alpha)\asymp N^{d-1},\quad f^{\prime}_{j_{0}}(\beta)\asymp(N^{\frac{d}{d_{0}}})^{d_{0}-1}=N^{d-\frac{d}{d_{0}}}.$
Dividing by $f^{\prime}_{1}(\alpha)$ in (2.15), we get
$E_{0}v-w\frac{f^{\prime}_{j_{0}}(\beta)}{f^{\prime}_{1}(\alpha)}=O\left(\frac{g(N)}{N^{d-1}}\right).$ (2.16)
Note that $E_{0}v$ is bounded and $g(N)/N^{d-1}\to 0$ $(N\to\infty)$ since $d\geq 2$. Also, we obtain
$\frac{f^{\prime}_{j_{0}}(\beta)}{f^{\prime}_{1}(\alpha)}\asymp\frac{N^{d-\frac{d}{d_{0}}}}{N^{d-1}}=N^{1-\frac{d}{d_{0}}}.$
If $d_{0}\geq d+1$, we get $f^{\prime}_{j_{0}}(\beta)/f^{\prime}_{1}(\alpha)\to\infty$ $(N\to\infty)$. This contradicts (2.16) since $w\geq 1$. Thus $d_{0}=d$, and hence $f^{\prime}_{j_{0}}(\beta)/f^{\prime}_{1}(\alpha)$ is bounded. Then by (2.16), we see that $w$ is a bounded positive integer.
Step 3. Let $f_{1}(x)\coloneqq\sum_{k=0}^{d}\lambda_{k}x^{k}$ and $f_{j_{0}}(x)\coloneqq\sum_{k=0}^{d}\psi_{k}x^{k}$ with some rational constants $\lambda_{k}$, $\psi_{k}$ $(k=0,1,\dots,d)$. In this step, we show
$\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}=\frac{w}{E_{0}v}.$ (2.17)
Since $\deg f_{1}(x)=\deg f_{j_{0}}(x)=d$ and $t,u\asymp N$, by (2.10) we obtain
$\lambda_{d}t^{d}-\psi_{d}u^{d}=O(N^{d-1})+O\bigl{(}g(N)\bigr{)}.$
Dividing by $t^{d}$, we get $\lambda_{d}-\psi_{d}(u/t)^{d}=O(1/N)$, and hence
$\biggl{(}\frac{u}{t}\biggr{)}^{d}=\frac{\lambda_{d}}{\psi_{d}}\left(1+O\Bigl{(}\frac{1}{N}\Bigr{)}\right).$
Here for every $\nu>0$, we have
$\left(1+O\Bigl{(}\frac{1}{N}\Bigr{)}\right)^{\nu}=1+O\Bigl{(}\frac{1}{N}\Bigr{)}.$ (2.18)
Then,
$\frac{u}{t}=\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}\left(1+O\Bigl{(}\frac{1}{N}\Bigr{)}\right).$ (2.19)
In addition, since $E_{0}v$ and $w$ are bounded, by (2.15) we obtain
$E_{0}vd\lambda_{d}t^{d-1}-wd\psi_{d}u^{d-1}=O(N^{d-2})+O\bigl{(}g(N)\bigr{)}.$
Dividing by $d\lambda_{d}t^{d-1}$, we get
$E_{0}v=w\frac{\psi_{d}}{\lambda_{d}}\biggl{(}\frac{u}{t}\biggr{)}^{d-1}+O\left(\frac{1}{N}\right)+O\left(\frac{g(N)}{N^{d-1}}\right).$
By (2.18) and (2.19), we obtain
$E_{0}v=w\frac{\psi_{d}}{\lambda_{d}}\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{d-1}{d}}\left(1+O\Bigl{(}\frac{1}{N}\Bigr{)}\right)+O\left(\frac{1}{N}\right)+O\left(\frac{g(N)}{N^{d-1}}\right)=w\left(\frac{\psi_{d}}{\lambda_{d}}\right)^{\frac{1}{d}}+O\left(\frac{g(N)}{N}\right).$
Since $E_{0}v$ is bounded, we obtain
$\frac{w}{E_{0}v}=\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}+O\left(\frac{g(N)}{N}\right).$
Note that $g(N)/N\to 0$ $(N\to\infty)$. Since $E_{0}v$ and $w$ are bounded integers, we get (2.17).
Step 4. Let $(a,b)$ be every pair with positive integers $a\in\mathcal{M}_{N}(j_{0})$ and $b$ satisfying
$|f_{1}(a)-f_{j_{0}}(b)|\leq g(N).\$ (2.20)
In this step, we show that there exists a rational number $\xi_{0}$ independent of $a$ and $b$ such that
$b=\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}\ a+\xi_{0}.$ (2.21)
Since $a\asymp N$, we get $b\asymp N$. Similarly to (2.19), by (2.20) we obtain
$\frac{b}{a}=\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}\left(1+O\Bigl{(}\frac{1}{N}\Bigr{)}\right).$
Multiplying by $a$, we get
$b=\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}a+O(1).$ (2.22)
Let
$\xi_{a,b}\coloneqq b-\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}a$ (2.23)
for each pair $(a,b)$ satisfying (2.20). By (2.22), we see that $\xi_{a,b}$ is bounded. Then, substituting $b=(\lambda_{d}/\psi_{d})^{1/d}\cdot a+\xi_{a,b}$ in (2.20), we obtain
$\lambda_{d}a^{d}+\lambda_{d-1}a^{d-1}-\psi_{d}\left\{\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}a\right\}^{d}-\psi_{d}d\left\{\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}a\right\}^{d-1}\xi_{a,b}-\psi_{d-1}\left\{\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}a\right\}^{d-1}$
=\ = $O(N^{d-2})+O\bigl{(}g(N)\bigr{)}.$
Dividing by $a^{d-1}$, we get
$\lambda_{d-1}-d\psi_{d}\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{d-1}{d}}\xi_{a,b}-\psi_{d-1}\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{d-1}{d}}=O\left(\frac{g(N)}{N}\right).$
Then, we get
$\xi_{a,b}=\frac{\lambda_{d-1}}{d\lambda_{d}}\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}-\frac{\psi_{d-1}}{d\psi_{d}}+O\left(\frac{g(N)}{N}\right).$
Note that $g(N)/N\to 0$ $(N\to\infty)$. Since $E_{0}v$ is a bounded integer, by (2.17) and (2.23), $\xi_{a,b}$ is a rational number which the denominator is bounded, and hence we obtain
$\xi_{a,b}=\frac{\lambda_{d-1}}{d\lambda_{d}}\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}-\frac{\psi_{d-1}}{d\psi_{d}}.$
Therefore, for each pair $(a,b)$ satisfying (2.20), $\xi_{a,b}$ is a rational number independent of $a$ and $b$. Thus by (2.23), we see that (2.21) holds.
Step 5. Finally, we show that (2.2) contradicts (2.1). By (2.9), (2.20) and (2.21), there exists an integer $D_{0}$ with $|D_{0}|\leq g(N)$ such that the following equation
$f_{1}(x)=f_{j_{0}}\left(\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}x+\xi_{0}\right)+D_{0}$
has at least $d+1$ solutions. Since $\deg f_{1}(x)=\deg f_{j_{0}}(x)=d$, we obtain
$f_{1}(x)\equiv f_{j_{0}}\left(\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}x+\xi_{0}\right)+D_{0}.$
Here by (2.10) and (2.21), we get
$u=\left(\frac{\lambda_{d}}{\psi_{d}}\right)^{\frac{1}{d}}t+\xi_{0}.$ (2.24)
Replacing $x$ with $x+t$, by (2.17) and (2.24), we obtain
$f_{1}(x+t)\equiv f_{j_{0}}\left(\frac{w}{E_{0}v}\ x+u\right)+D_{0}.$ (2.25)
Now comparing the leading coefficients, we get $r_{1}/s_{1}=(r_{j_{0}}/s_{j_{0}})\{w/(E_{0}v)\}^{d}$, namely,
$r_{1}s_{j_{0}}E_{0}^{d}=r_{j_{0}}s_{1}\left(\frac{w}{v}\right)^{d}.$
By (2.3), we get $r_{j_{0}}s_{1}\mid E_{0}$. Then, $B_{0}\coloneqq w/v$ is a positive integer since $v$ and $w$ are positive integers. Also since $t\in\mathcal{M}_{N}(j_{0})$, there exists a positive integer $T$ such that $t=E_{0}T+A$. By (2.25), replacing $x$ with $x-E_{0}T$, we obtain
$f_{1}(x+A)\equiv f_{j_{0}}\left(\frac{B_{0}}{E_{0}}x+(u-B_{0}T)\right)+D_{0}.$
This contradicts (2.1). Therefore, the proof of Lemma 2.1 is completed. ∎
3 Proof of Theorem 1.1
In this section, we prove Theorem 1.1.
Proof.
Suppose to the contrary that the numbers (1.1) in Theorem 1.1 are linearly dependent over $\mathbb{Q}(q)$. Then, there exist algebraic integers $b_{j}$ $(j=0,1,\dots,\ell)$ in $\mathbb{Q}(q)$ not all zero such that
$b_{0}+\sum_{j=1}^{\ell}b_{j}\sum_{n=1}^{\infty}\frac{a_{j}(n)}{q^{f_{j}(n)}}=0.$ (3.1)
Let
$\mathcal{R}$$\coloneqq\{j\in\mathbb{Z}_{\geq 1}\mid b_{j}\neq 0,\ j=1,2,\dots,\ell\},$
$\mathcal{S}$$\coloneqq\bigl{\{}j\in\mathcal{R}\mid\deg f_{j}(x)=\min_{k\in\mathcal{R}}\{\deg f_{k}(x)\}\bigr{\}}.$
Since $\mathcal{R}\neq\emptyset$, we also obtain $\mathcal{S}\neq\emptyset$. By the conditions (i), (ii) in Theorem 1.1, for the non-empty subset $\mathcal{S}$ of $\{1,2,\dots,\ell\}$, there exist integers $i\in\mathcal{S}$ and $A$ satisfying the following two conditions:
  • (I)
    $f_{i}(x+A)\not\equiv f_{j}(Bx+C)+D$ for any integers $j\in\mathcal{S}\setminus\{i\}$, $C$, $D$ and any positive rational number $B$.
  • (II)
    There exists a positive integer $E$ such that $\liminf_{n\to\infty}|a_{i}(En+A)|\neq 0$.
Without loss of generality, we may assume $i=1$. Then, we get
$f_{1}(x+A)\not\equiv f_{j}(Bx+C)+D$ (3.2)
for any integers $j\in\mathcal{R}\setminus\{1\}$, $C$, $D$ and any positive rational number $B$, and
$\liminf_{n\to\infty}|a_{1}(En+A)|\neq 0.$ (3.3)
Next, we apply Lemma 2.1. We define
$h(x)\coloneqq 1+\max_{\begin{subarray}{c}n\\ x\geq n\in\mathbb{Z}_{\geq 1}\end{subarray}}\left\{\sum_{j\in\mathcal{R}}\sum_{\sigma}\log_{q}(1+|a_{j}(n)^{\sigma}|)\right\}\quad(x\geq 1),$ (3.4)
where $\sigma$ runs through all embeddings of $\mathbb{Q}(q)\to\mathbb{C}$. Also, we define
$H(x)$ $\coloneqq\min_{\begin{subarray}{c}y\\ x\leq y\end{subarray}}\frac{y}{h(y)}\quad(x\geq 1),$ (3.5)
$G(x)$ $\coloneqq x\cdot H(x)^{-\frac{1}{4}}\quad(x\geq 1).$ (3.6)
By (3.4), we obtain $h(x)=o(x)$ since $\log(1+|a_{j}(n)^{\sigma}|)=o(n)$ for every $j\in\mathcal{R}$ and every embedding $\sigma\colon\mathbb{Q}(q)\to\mathbb{C}$. Then by (3.5), we get
$H(x)\to\infty\quad(x\to\infty),$ (3.7)
and hence by (3.6), we obtain
$G(x)=o(x).$ (3.8)
Also by (3.4), we get $h(x)\geq 1$ for $x\geq 1$. Then by (3.5), we get
$0<H(x)\leq\frac{x}{h(x)}\leq x\quad(x\geq 1),$ (3.9)
and hence by (3.6), we obtain
$G(x)\geq x^{\frac{3}{4}}\to\infty\quad(x\to\infty).$ (3.10)
By (3.2), (3.8) and (3.10), Lemma 2.1 implies that there exist infinitely many positive integers $m$ with $m\equiv A\pmod{E}$ such that
$|f_{1}(m)-f_{j}(k)|>G(m)$ (3.11)
for any positive integers $j\in\mathcal{R}$ and $k$ with $(j,k)\neq(1,m)$.
Let $m$ be a sufficiently large integer satisfying (3.11). By the hypotheses, note that $f_{j}(x)$ $(j\in\mathcal{R})$ are increasing for large $x$. We define
$n_{m,j}$ $\coloneqq\max\{n\in\mathbb{Z}_{\geq 1}\mid f_{j}(n)<f_{1}(m)-G(m)\}\quad(j\in\mathcal{R}),$ (3.12)
$N_{m,j}$ $\coloneqq\min\{n\in\mathbb{Z}_{\geq 1}\mid f_{j}(n)>f_{1}(m)+G(m)\}\quad(j\in\mathcal{R}),$ (3.13)
$K_{m}$ $\coloneqq\max_{j\in\mathcal{R}}\{f_{j}(n_{m,j})\}.$ (3.14)
Multiplying (3.1) by $q^{K_{m}}$, by (3.11), (3.12) and (3.13), we obtain
$b_{0}q^{K_{m}}+\sum_{j\in\mathcal{R}}b_{j}\sum_{n=1}^{n_{m,j}}a_{j}(n)q^{K_{m}-f_{j}(n)}+b_{1}\ \frac{a_{1}(m)}{q^{f_{1}(m)-K_{m}}}+\sum_{j\in\mathcal{R}}b_{j}\sum_{n=N_{m,j}}^{\infty}\frac{a_{j}(n)}{q^{f_{j}(n)-K_{m}}}=0.$
Let
$P_{m}(X)\coloneqq-b_{0}X^{K_{m}}-\sum_{j\in\mathcal{R}}b_{j}\sum_{n=1}^{n_{m,j}}a_{j}(n)X^{K_{m}-f_{j}(n)}.$ (3.15)
Then,
$P_{m}(q)=\frac{a_{1}(m)}{q^{f_{1}(m)-K_{m}}}\left\{b_{1}+\sum_{j\in\mathcal{R}}b_{j}\sum_{n=0}^{\infty}\frac{a_{j}(N_{m,j}+n)}{a_{1}(m)q^{f_{j}(N_{m,j}+n)-f_{1}(m)}}\right\}.$ (3.16)
Now, we evaluate (3.16). By $\deg f_{1}(x)\leq\deg f_{j}(x)$ $(j\in\mathcal{R})$ and (3.8), there exists a positive integer $c_{0}$ independent of $m$ such that $f_{j}(c_{0}m)>f_{1}(m)+G(m)$ $(j\in\mathcal{R})$. Then by (3.7), (3.12) and (3.13), we get
$n_{m,j}<N_{m,j}\leq c_{0}m<m\cdot H(m)^{\frac{1}{2}}\quad(j\in\mathcal{R}).$ (3.17)
By (3.9), we have $h(x)\leq x/H(x)$ $(x\geq 1)$. In addition, by (3.4) and (3.5), $h(x)$ and $H(x)$ $(x\geq 1)$ are increasing functions. Then by (3.7) and (3.17), for every integers $j\in\mathcal{R}$ and $n\geq 0$, we obtain
$h(N_{m,j}+n)\leq h\bigl{(}m\cdot H(m)^{\frac{1}{2}}+n\bigr{)}\leq\frac{m\cdot H(m)^{\frac{1}{2}}+n}{H\bigl{(}m\cdot H(m)^{\frac{1}{2}}+n\bigr{)}}\leq\frac{m\cdot H(m)^{\frac{1}{2}}+n}{H(m)}\leq m\cdot H(m)^{-\frac{1}{2}}+n.$
Then by (3.6), for every integers $j\in\mathcal{R}$ and $n\geq 0$, we obtain
$h(N_{m,j}+n)\leq G(m)\cdot H(m)^{-\frac{1}{4}}+n.$ (3.18)
Let $c_{1},c_{2},\dots$ be some positive constants independent of $m$. By (3.3), we get $|a_{1}(m)|>c_{1}$. Also by (3.4), for every $j\in\mathcal{R}$ and every embedding $\sigma\colon\mathbb{Q}(q)\to\mathbb{C}$, we obtain
$|a_{j}(n)^{\sigma}|\leq q^{h(n)}\quad(n\geq 1).$
Since $h(x)$ $(x\geq 1)$ is an increasing function, by (3.13), we get $h(m)\leq h(N_{m,1})$. In addition, note that $G(x)\to\infty$ and $H(x)\to\infty$ $(x\to\infty)$. By (3.12), (3.14) and (3.18), we get
$0<\left|\frac{a_{1}(m)}{q^{f_{1}(m)-K_{m}}}\right|\leq\frac{q^{h(m)}}{q^{G(m)}}\leq\frac{q^{h(N_{m,1})}}{q^{G(m)}}\leq q^{-G(m)\left(1-H(m)^{-1/4}\right)}\to 0\quad(m\to\infty).$ (3.19)
Here by (3.13), if $m$ is large, $N_{m,j}$ is also large for every $j\in\mathcal{R}$. Then, $f_{j}(x+1)-f_{j}(x)\geq 2$ for $x\geq N_{m,j}$ $(j\in\mathcal{R})$. Thus by (3.13) and (3.18), for each $j\in\mathcal{R}$, we obtain
$\left|\sum_{n=0}^{\infty}\frac{a_{j}(N_{m,j}+n)}{a_{1}(m)q^{f_{j}(N_{m,j}+n)-f_{1}(m)}}\right|$ $\leq c_{2}\ \frac{1}{q^{f_{j}(N_{m,j})-f_{1}(m)}}\sum_{n=0}^{\infty}\frac{q^{h(N_{m,j}+n)}}{q^{f_{j}(N_{m,j}+n)-f_{j}(N_{m,j})}}$
$\leq c_{2}\ \frac{q^{G(m)\cdot H(m)^{-1/4}}}{q^{G(m)}}\sum_{n=0}^{\infty}\frac{q^{n}}{q^{2n}}$
$=c_{2}\ \frac{q}{q-1}\ q^{-G(m)\left(1-H(m)^{-1/4}\right)}\to 0\quad(m\to\infty).$ (3.20)
In addition, note that $b_{j}$ $(j\in\mathcal{R})$ are non-zero constants of algebraic integers in $\mathbb{Q}(q)$. By (3.16), (3.19) and (3.20), we obtain
$0<|P_{m}(q)|\leq c_{3}q^{-G(m)\left(1-H(m)^{-1/4}\right)}\to 0\quad(m\to\infty).$ (3.21)
Finally, we evaluate the norm over $\mathbb{Q}(q)/\mathbb{Q}$ of the number $P_{m}(q)$. Note that $h(x)$ $(x\geq 1)$ is an increasing function. Then, for every embedding $\sigma\colon\mathbb{Q}(q)\to\mathbb{C}$ with $q^{\sigma}\neq q$, by $|q^{\sigma}|\leq 1$, (3.14), (3.15), (3.17) and (3.18), we get
$|P_{m}(q)^{\sigma}|$ $\leq|b_{0}^{\sigma}|\cdot|q^{\sigma}|^{K_{m}}+\sum_{j\in\mathcal{R}}|b_{j}^{\sigma}|\sum_{n=1}^{n_{m,j}}|a_{j}(n)^{\sigma}|\cdot|q^{\sigma}|^{K_{m}-f_{j}(n)}$
$\leq c_{4}\sum_{j\in\mathcal{R}}\sum_{n=1}^{n_{m,j}}|a_{j}(n)^{\sigma}|\leq c_{4}\sum_{j\in\mathcal{R}}n_{m,j}q^{h(N_{m,j})}\leq c_{5}mq^{G(m)\cdot H(m)^{-1/4}}.$ (3.22)
By (3.14), (3.15) and (3.21), we see that $P_{m}(q)$ is a non-zero algebraic integer in $\mathbb{Q}(q)$. Thus by (3.7), (3.10), (3.21) and (3.22), we obtain
$1\leq|N_{\mathbb{Q}(q)/\mathbb{Q}}\ P_{m}(q)|$ $\leq c_{3}q^{-G(m)\left(1-H(m)^{-1/4}\right)}\left(c_{5}mq^{G(m)\cdot H(m)^{-1/4}}\right)^{[\mathbb{Q}(q):\mathbb{Q}]-1}$
$\leq c_{6}m^{[\mathbb{Q}(q):\mathbb{Q}]-1}q^{-m^{3/4}\left(1-[\mathbb{Q}(q):\mathbb{Q}]\cdot H(m)^{-1/4}\right)}\to 0\quad(m\to\infty).$
This is a contradiction for the sufficiently large integer $m$. Therefore, the proof of Theorem 1.1 is completed. ∎
4 Proof of Theorem 1.2
In this section, we prove Theorem 1.2.
Proof.
Suppose that there exist algebraic integers $b_{0}$, $b_{i,j}$ $\bigl{(}(i,j)\in\mathcal{U}\bigr{)}$ in $\mathbb{Q}(q)$ such that
$b_{0}+\sum_{i=1}^{m}\sum_{j=1}^{\ell_{i}}b_{i,j}\sum_{n=1}^{\infty}\frac{P_{i,j}(n)}{q^{f_{i,j}(n)}}=0.$
For each $(i,j)\in\mathcal{U}$, we get
$\sum_{n=1}^{\infty}\frac{P_{i,j}(n)}{q^{f_{i,j}(n)}}=\sum_{n=0}^{\infty}\frac{\tilde{p}_{i,j}(n)}{q^{f_{i,j}\bigl{(}\frac{n+C_{i,j}}{B_{i,j}}\bigr{)}}}+s_{i,j}=\sum_{n=0}^{\infty}\frac{q^{D_{i,j}}\cdot\tilde{p}_{i,j}(n)}{q^{g_{i}(n)}}+s_{i,j},$
where $s_{i,j}\in\mathbb{Q}(q)$ is some constant. Let $d$ be the minimum of $D_{i,j}$ $\bigl{(}(i,j)\in\mathcal{U}\bigr{)}$. Also, for each $i=1,2,\dots,m$, let $K_{i}$ be the least common multiple of $B_{i,j}$ $(j=1,2,\dots,\ell_{i})$. Then,
$s_{0}+q^{d}\sum_{i=1}^{m}\sum_{r=0}^{K_{i}-1}\sum_{n=0}^{\infty}\frac{\displaystyle\sum_{j=1}^{\ell_{i}}b_{i,j}q^{D_{i,j}-d}\cdot\tilde{p}_{i,j}(K_{i}n+r)}{q^{g_{i}(K_{i}n+r)}}=0,$
where $s_{0}\coloneqq b_{0}+\sum_{(i,j)\in\mathcal{U}}b_{i,j}s_{i,j}\in\mathbb{Q}(q)$. Note that $K_{i}/B_{i,j}$ $\bigl{(}(i,j)\in\mathcal{U}\bigr{)}$ are positive integers. For each $(i,j)\in\mathcal{U}$, since
[…]
we see that for each $i=1,2,\dots,m$ and $r=0,1,\dots,K_{i}-1$,
$\sum_{j=1}^{\ell_{i}}b_{i,j}q^{D_{i,j}-d}\cdot\tilde{p}_{i,j}(K_{i}n+r)$ (4.2)
is a polynomial of $n$ with algebraic integer coefficients in $\mathbb{Q}(q)$. Now, we define
[…]
We show $\mathcal{V}=\emptyset$.
Suppose to the contrary $\mathcal{V}\neq\emptyset$. For each $(i,r)\in\mathcal{V}$, let $h_{i,r}(x)\coloneqq g_{i}(K_{i}x+r)$. Then,
$s_{0}+q^{d}\sum_{(i,r)\in\mathcal{V}}\sum_{n=0}^{\infty}\frac{\displaystyle\sum_{j=1}^{\ell_{i}}b_{i,j}q^{D_{i,j}-d}\cdot\tilde{p}_{i,j}(K_{i}n+r)}{q^{h_{i,r}(n)}}=0.$ (4.3)
We apply Theorem 1.1. By the definition of $\mathcal{V}$, we obtain
$\liminf_{n\to\infty}\left|\sum_{j=1}^{\ell_{i}}b_{i,j}q^{D_{i,j}-d}\cdot\tilde{p}_{i,j}(K_{i}n+r)\right|\neq 0\quad\bigl{(}(i,r)\in\mathcal{V}\bigr{)}.$
By (4.1), if $(i,r)\in\mathcal{V}$, there exists an integer $j_{0}$ $(1\leq j_{0}\leq\ell_{i})$ such that $(r+C_{i,j_{0}})/B_{i,j_{0}}$ is an integer, and hence
$h_{i,r}(x)\equiv f_{i,j_{0}}\left(\frac{K_{i}}{B_{i,j_{0}}}x+\frac{r+C_{i,j_{0}}}{B_{i,j_{0}}}\right)+D_{i,j_{0}}$
is an integer-valued polynomial of degree $\geq 2$ with positive leading coefficients. Also, for each $i=1,2,\dots,m$, we obtain
$h_{i,r_{1}}(x+A)\not\equiv h_{i,r_{2}}(Bx+C)+D$ (4.4)
for any integers $10\leq r_{1},\ r_{2}\leq K_{i}-$ 1 $(r_{1}\neq r_{2})$, $A$, $C$, $D$ and any positive rational number $B$. Indeed, suppose to the contrary that there exist integers $i$ $(1\leq i\leq m)$, $r_{1}$, $r_{2}$ $(0\leq r_{1},\ r_{2}\leq K_{i}-1,$ $r_{1}\neq r_{2})$, $A$, $C$, $D$ and a positive rational number $B$ such that
$h_{i,r_{1}}(x+A)\equiv h_{i,r_{2}}(Bx+C)+D.$
Namely,
$g_{i}(K_{i}x+AK_{i}+r_{1})\equiv g_{i}(BK_{i}x+CK_{i}+r_{2})+D.$
Then, we get $B=1$, $AK_{i}+r_{1}=CK_{i}+r_{2}$ and $D=0$. In particular, we get $r_{1}-r_{2}=K_{i}(C-A)$, and this contradicts $10\leq r_{1},\ r_{2}\leq K_{i}-$ 1 and $r_{1}\neq r_{2}$. Thus, we see that (4.4) holds. In addition, we have
$g_{i_{1}}(x)\not\sim g_{i_{2}}(x)\quad(1\leq i_{1}<i_{2}\leq m).$
Hence, we obtain
$h_{i_{1},r_{1}}(x+A)\not\equiv h_{i_{2},r_{2}}(Bx+C)+D$
for any $(i_{1},r_{1})\neq(i_{2},r_{2})$, any integers $A$, $C$, $D$ and any positive rational number $B$. Therefore, by Theorem 1.1, the numbers
$1,\qquad\sum_{n=0}^{\infty}\frac{\displaystyle\sum_{j=1}^{\ell_{i}}b_{i,j}q^{D_{i,j}-d}\cdot\tilde{p}_{i,j}(K_{i}n+r)}{q^{h_{i,r}(n)}}\quad\bigl{(}(i,r)\in\mathcal{V}\bigr{)}$
are linearly independent over $\mathbb{Q}(q)$, and this contradicts (4.3). Thus $\mathcal{V}=\emptyset$, and hence for each $i=1,2,\dots,m$, we obtain
$\sum_{j=1}^{\ell_{i}}b_{i,j}q^{D_{i,j}-d}\cdot\tilde{p}_{i,j}(K_{i}n+r)=0$
for every integers $10\leq r\leq K_{i}-$ 1 and $n\geq 0$. Therefore, the numbers (1.7) in Theorem 1.2 are linearly independent over $\mathbb{Q}(q)$ if and only if for each $i=1,2,\dots,m$, the sequences $\{\tilde{p}_{i,j}(n)\}_{n\geq 0}$ $(j=1,2,\dots,\ell_{i})$ are linearly independent over $\mathbb{Q}(q)$. The proof of Theorem 1.2 is completed. ∎
5 Examples
5.1 Examples of Theorem 1.1
First, we give some examples of the set of polynomials which satisfy the condition (i) in Theorem 1.1. We give the following lemma to prove Example 1.2 in Section 1.
Lemma 5.1.
Let $g(x)$ be any integer-valued polynomial of degree $d\geq 1$ with a positive leading coefficient. Let
$f_{i,j}(x)\coloneqq ig(x)^{j}+h_{i,j}(x)\quad\bigl{(}i=1,2,\dots,\ell_{1},\ j=1,2,\dots,\ell_{2}\text{ with }j\geq\frac{2}{d}\bigr{)},$ (5.1)
where $h_{i,j}(x)$ $(i=1,2,\dots,\ell_{1},\ j=1,2,\dots,\ell_{2}\text{ with }j\geq 2/d)$ are any integer-valued polynomials of degree $\leq dj-2$. Then, the polynomials (5.1) satisfy the condition (i) in Theorem 1.1.
Proof.
Let
$A_{0}\coloneqq-\left\lfloor\frac{u}{td}\right\rfloor-1,$
where $t>0$ and $u$ are coefficients of $x^{d}$ and $x^{d-1}$ in $g(x)$, respectively. Let $\mathcal{S}$ be any non-empty subset of $\{(i,j)\mid i=1,2,\dots,\ell_{1},\ j=1,2,\dots,\ell_{2}\text{ with }j\geq 2/d\}$ and $(i_{1},j_{1})\in\mathcal{S}$ be an element satisfying $i_{1}=\min\{i\mid(i,j)\in\mathcal{S}\}$. Suppose to the contrary that there exist $(i_{2},j_{2})\in\mathcal{S}\setminus\{(i_{1},j_{1})\}$, a positive rational number $B$ and integers $C$, $D$ such that
$f_{i_{1},j_{1}}(x+A_{0})\equiv f_{i_{2},j_{2}}(Bx+C)+D.$
Namely,
$i_{1}g(x+A_{0})^{j_{1}}+h_{i_{1},j_{1}}(x+A_{0})\equiv i_{2}g(Bx+C)^{j_{2}}+h_{i_{2},j_{2}}(Bx+C)+D.$
Then, we get $j_{1}=j_{2}$, and hence $i_{1}<i_{2}$. By comparing the coefficients of $x^{dj_{1}}$ and $x^{dj_{1}-1}$, we obtain
$i_{1}t^{j_{1}}=i_{2}(tB^{d})^{j_{1}},$ (5.2)
$i_{1}j_{1}t^{j_{1}-1}(tdA_{0}+u)=i_{2}j_{1}(tB^{d})^{j_{1}-1}B^{d-1}(tdC+u),$ (5.3)
respectively. By (5.2), we obtain $i_{1}=i_{2}(B^{d})^{j_{1}}$. Since $i_{1}<i_{2}$, we get $10<B<$ 1. In addition, substituting in (5.3), we obtain $B(tdA_{0}+u)=tdC+u$. Then,
$C$ $=BA_{0}-(1-B)\frac{u}{td}$
$=B\left(-\left\lfloor\frac{u}{td}\right\rfloor-1\right)-(1-B)\left(\left\lfloor\frac{u}{td}\right\rfloor+\left\{\frac{u}{td}\right\}\right)$
$=-\left\lfloor\frac{u}{td}\right\rfloor-1+(1-B)\left(1-\left\{\frac{u}{td}\right\}\right),$
where $\{x\}\coloneqq x-\lfloor x\rfloor$ denotes the fractional part of $x$. Note that 0<($10<(1-B)\bigl{(}1-\{u/(td)\}\bigr{)}<$ 1. This is a contradiction since $C$ is an integer. Thus, we obtain
$f_{i_{1},j_{1}}(x+A_{0})\not\equiv f_{i_{2},j_{2}}(Bx+C)+D$
for any $(i_{2},j_{2})\in\mathcal{S}\setminus\{(i_{1},j_{1})\}$, any positive rational number $B$ and any integers $C$, $D$. Therefore, the proof of Lemma 5.1 is completed. ∎
By Corollary 1.1 in Section 1, we get the following corollary.
Corollary 5.1.
Let $q$ be a Pisot or Salem number. Let $f_{i,j}(x)$ $(i=1,2,\dots,\ell_{1},\ j=1,2,\dots,\ell_{2}$ with $j\geq 2/d)$ be polynomials defined by (5.1) in Lemma 5.1. Then, for any sequences $\{a_{i,j}(n)\}_{n\geq 1}$ $(i\!=1,2,\dots,\ell_{1},\ j\!=1,2,\dots,\ell_{2}\text{ with }j\geq 2/d)$ of non-zero integers with $\log(1+|a_{i,j}(n)|)=o(n)$, the numbers
$1,\qquad\sum_{n=1}^{\infty}\frac{a_{i,j}(n)}{q^{f_{i,j}(n)}}\quad\bigl{(}i=1,2,\dots,\ell_{1},\ j=1,2,\dots,\ell_{2}\text{ with }j\geq\frac{2}{d}\bigr{)}$
are linearly independent over $\mathbb{Q}(q)$.
In Lemma 5.1, if $g(x)\equiv x$, $h_{i,j}(x)\equiv 0$ and $\ell_{1}=\ell_{2}=\ell$, we get Example 1.2 in Section 1. Also in Lemma 5.1, if $g(x)\equiv x(x+1)$ or $x^{2}$, $h_{i,j}(x)\equiv 0$, $\ell_{1}=\ell$ and $\ell_{2}=1$, we get linear independence results for the values of the Jacobi theta constants defined by (1.2) in Section 1.
Example 5.1.
Let $q$ be a Pisot or Salem number. Then, for any integer $\ell\geq 1$, the numbers
$1,\qquad q^{\frac{i}{4}}\theta_{2}(1/q^{i})\quad(i=1,2,\dots,\ell)$
are linearly independent over $\mathbb{Q}(q)$. So are the numbers
$1,\qquad\theta_{3}(1/q^{i})\quad(i=1,2,\dots,\ell),$
and the numbers
$1,\qquad\theta_{4}(1/q^{i})\quad(i=1,2,\dots,\ell).$
Note that the numbers $1$, $q^{i/4}\theta_{2}(1/q^{i})$, $\theta_{3}(1/q^{i})$ $(i=1,2,\dots,\ell)$ are linearly dependent over $\mathbb{Q}(q)$ if $\ell\geq 4$. For example, $\theta_{3}(1/q)=\theta_{2}(1/q^{4})+\theta_{3}(1/q^{4})$ since
$1+2\sum_{n=1}^{\infty}\frac{1}{q^{n^{2}}}=2\sum_{n=0}^{\infty}\frac{1}{q^{(2n+1)^{2}}}+\left(1+2\sum_{n=1}^{\infty}\frac{1}{q^{(2n)^{2}}}\right).$
Also, it is known that for any integer $\ell\geq 1$ and any algebraic number $\alpha$ $(0<|\alpha|<1)$, any two numbers among the numbers $\theta_{m}(\alpha^{i})$ $(m=2,3,4,\ i=1,2,\dots,\ell)$ are algebraically independent over $\mathbb{Q}$, while any three are not (cf. [3, 4, 5]).
5.2 An example of Theorem 1.2
Next, we give the following example which Theorem 1.2 is effective.
Example 5.2.
Let $q$ be a Pisot or Salem number. Let $P_{1}(x)$, $P_{2}(x)$ and $P_{3}(x)$ be non-zero polynomials with algebraic integer coefficients in $\mathbb{Q}(q)$. Then, the four numbers
$1,\quad\sum_{n=1}^{\infty}\frac{P_{1}(n)}{q^{n^{2}}},\quad\sum_{n=1}^{\infty}\frac{P_{2}(n)}{q^{(2n-1)^{2}}},\quad\sum_{n=1}^{\infty}\frac{P_{3}(n)}{q^{(2n)^{2}}}$
are linearly dependent over $\mathbb{Q}(q)$ if and only if there exist algebraic numbers $c_{1}\neq 0$ and $c_{2}\neq 0$ in $\mathbb{Q}(q)$ such that
$P_{2}(x)\equiv c_{1}P_{1}(2x-1),\quad P_{3}(x)\equiv c_{2}P_{1}(2x).$ (5.4)
Proof.
We apply the criterion given by Theorem 1.2. Let $f_{1}(x)\coloneqq x^{2}$, $f_{2}(x)\coloneqq(2x-1)^{2}$, $f_{3}(x)\coloneqq(2x)^{2}$ and $g(x)\coloneqq x^{2}$. Then, we get $f_{1}\sim f_{2}\sim f_{3}$ and
$g(x)\equiv f_{1}(x)\equiv f_{2}\Bigl{(}\frac{x+1}{2}\Bigr{)}\equiv f_{3}\Bigl{(}\frac{x}{2}\Bigr{)}.$
Let $\{\tilde{p}_{j}(n)\}_{n\geq 0}$ $(j=1,2,3)$ be sequences such that
$\tilde{p}_{1}(n)$ $\coloneqq P_{1}(n)\quad\text{for every integer $n$,}$
[…]
Assume that there exist algebraic numbers $k_{1}$, $k_{2}$ and $k_{3}$ in $\mathbb{Q}(q)$ such that
$k_{1}\cdot\tilde{p}_{1}(n)+k_{2}\cdot\tilde{p}_{2}(n)+k_{3}\cdot\tilde{p}_{3}(n)=0$
for every integer $n\geq 0$. Then, we obtain
[…]
Thus, we see that the sequences $\{\tilde{p}_{j}(n)\}_{n\geq 0}$ $(j=1,2,3)$ are linearly dependent over $\mathbb{Q}(q)$ if and only if the condition (5.4) holds. Therefore, by Theorem 1.2, the proof of Example 5.2 is completed. ∎
Acknowledgments. I would like to express my sincere gratitude to Y. Tachiya for his comments and suggestions on this manuscript.